OpenStudy (anonymous):

Simplify the expression. 2 over 3(-9m+12) A. -6m + 12 B. -18m + 8 C. -6m + 24 D. -6m + 8

OpenStudy (unklerhaukus):

is that \[{2 \over 3}(-9m+12)\] or \[{2 \over 3(-9m+12)}\]?

OpenStudy (anonymous):

First one

OpenStudy (unklerhaukus):

ok so distribute that fraction across the brackets \[\frac23(−9m+12)=\left(\frac23\times-9m\right)+\left(\frac23\times12\right)\]

OpenStudy (unklerhaukus):

now one of the factors of nine , is three, so that will cancel, likewise for the twelve,,